# Redefining non-IID Data in Federated Learning for Computer Vision Tasks: Migrating from Labels to Embeddings for Task-Specific Data Distributions

### v2

@@ Abstract
Federated Learning (FL) represents a paradigm shift in distributed machine learning (ML), enabling clients to train models collaboratively while keeping their raw data private. This paradigm shift from traditional centralized ML introduces challenges due to the non-iid (non-independent and identically distributed) nature of data across clients, significantly impacting FL’s performance. Existing literature predominantly model data heterogeneity by imposing label distribution skew across clients. In this paper, we show that label distribution skew fails to fully capture the real-world data heterogeneity among clients in computer vision tasks beyond classification. Subsequently, we demonstrate that current approaches overestimate FL’s performance by relying on label/class distribution skew, exposing an overlooked gap in the literature. By utilizing pre-trained deep neural networks to extract task-specific data embeddings, we define task-specific data heterogeneity through the lens of each vision task and introduce a new level of data heterogeneity called embedding-based data heterogeneity. Our methodology involves clustering data points based on embeddings and distributing them among clients using the Dirichlet distribution. Through extensive experiments, we evaluate the performance of different FL methods under our revamped notion of data heterogeneity, introducing new benchmark performance measures to the literature. We further unveil a series of open research directions that can be pursued.
@@ I-A Overview and Literature Review
To bridge this gap, we redefine the notion of data heterogeneity in FL for computer vision tasks as follows:
@@
To achieve the above objective, for the first time in the literature, we introduce a new layer of data heterogeneity, which we call embedding-based data heterogeneity. To model embedding-based data heterogeneity (i.e., embedding distribution skew), we leverage a pre-trained deep neural network (DNN) for each task and feed forward data points through this network. We then exploit the fact that the initial layers of the DNN of each task extract low-level features, and as the data progresses through deeper layers, the network extracts higher-level representations that are increasingly specific to the vision task of interest. Particularly, as we reach the layer just prior to the last layer (i.e, penultimate layer), embeddings that encapsulate a comprehensive view of the data from the perspective of the task are obtained, integrating both low-level and high-level features. We then leverage K-means to cluster datapoints based on these embeddings. Subsequently, we utilize a Dirichlet distribution to assign data points of each cluster to different clients to emulate task-specific data heterogeneity.
@@ I-B Summary of Contributions
We highlight the significance of task similarity in FL, illustrating that differing tasks may interpret the same data distribution across clients with varying degrees of heterogeneity. Through this nuanced observation, we unveil a series of novel research directions on addressing data heterogeneity for vision tasks in single- and multi-task FL settings.
@@ II-B4 FL for Non-Classification Vision Tasks under Class-based Data Heterogeneity
We next take one of the well-known computer vision datasets, called Taskonomy [15], due to the large number of tasks it includes which is why it has been previously used in both centralized [23] and federated [16] multi-task learning as a sole baseline. We utilize the above-described Dirichlet distribution to create a heterogeneous allocation of data across $25$ clients’ local datasets in FL using the available class labels, referred to as “class_scene” in this dataset. [^4] The classes are provided in the dataset as probabilities for each datapoint belonging to each available class from which we choose the class with the highest probability as the determined class for each datapoint. More details on the data processing is provided in Appendix A. In our experiments, we select a representative set of vision tasks in this dataset, which are visualized in Fig. 2 and include “2D Edges”, “Reshading”, “Surface Normals”, “Semantic Segmentation”, “3D Keypoints”, “Euclidean Depth Estimation”, and “Scene Classification.”
These vision tasks – except Scene Classification – can be construed as non-classification tasks since they are not associated with a class label. Following the non-iid data induced by Dirichlet distribution across the clients, we study the performance of the FL global model under different levels of heterogeneity – various $\alpha$ values – for the aforementioned vision tasks. The results of this experiment are shown in Fig. 3 (the left plot in each box, titled “Class-based”).
@@ III-B Emulating Heterogeneous Local Datasets for Clients
Formally, once the clustering of datapoints based on their embeddings are obtained, we emulate the local datasets $\mathcal{D}_{n}$ for each client $n$, where the set of ratios for classes $\mathbf{p}$ detailed in Sec. II-B3 is now replaced by the distribution of the embeddings in each cluster/group, which we denote by $\widehat{\mathbf{p}}$. Also, we let the set $\mathcal{G}$ contain the distinct cluster/group indices (e.g., all $16$ groups in Fig. 5). In our approach, each client $n$ will be assigned data from each cluster/group based on its own cluster/group distribution vector $\widehat{\mathbf{q}}_{n}=\mathsf{Dir}(\alpha\widehat{\mathbf{p}})=\{\widehat{q}_{n,g}\}_{g\in\mathcal{G}}$ where $\widehat{q}_{n,g}$ denotes the ratio of client $n$’s dataset containing data from group $g$ (e.g., one of the $16$ groups in Fig. 5). These local datasets are then used during the local training of FL.
@@
Remark 2 (Implications on the Concept of Task Similarity).
@@
Following the footsteps of prior work on modeling task similarity in the centralized ML domain [28, 29], our approach opens a model-driven perspective to quantify the similarity between two tasks. Specifically, this involves clustering data points based on their embeddings obtained from the models pre-trained for each task and comparing the resulting clusters to assess similarity. Using this straightforward approach, we establish a baseline performance for task similarity analysis, with results discussed in Sec. IV-B2. While this serves as a preliminary exploration, more sophisticated methods for task similarity measurement in the FL setting remain an open area for future investigation. Furthermore, this analysis provides an avenue for developing new task grouping strategies in FL, akin to prior efforts in centralized ML [30]. For example, by employing the task similarity measures based on the embedding-based clustering of data points, tasks with highly similar clustering patterns can be grouped and trained collaboratively, potentially enhancing FL performance for related tasks. We leave further exploration of this notion as future work.
Remark 3 (Implications on Multi-Task FL and its Extensions).
@@
A fundamental implications of our method lies in its application to multi-task FL and its extensions. In conventional multi-task FL [19, 18], each client typically has a single local task, with tasks varying across clients. On the other hand, the extended variations of multi-task FL [16, 17] consider scenarios where multiple tasks coexist on a single FL client. In the context of conventional multi-task FL, our method enables researchers to explore how varying levels of data heterogeneity, from the perspectives of different tasks, affect the performance of multi-task FL systems. This opens the door to designing countermeasures specifically aimed at addressing task-induced data heterogeneity in multi-task FL. Similarly, for the extensions of multi-task FL, our method allows researchers to investigate how inducing data heterogeneity from the perspective of one task impacts the performance of other tasks trained concurrently on the same client. Moreover, it facilitates the study of grouping and training local tasks together under different levels of data heterogeneity, potentially leading to new strategies for optimizing task interactions in FL. We leave further exploration of these concepts to future work.
@@
IV Experiments and Benchmarks
IV-A FL Methods
  1. 1.
    FedAvg[2]: One of the widely used methods in FL with the dynamics explained in Sec. II-A. The results presented in Fig. 3 were also gathered by employing this method.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
@@ IV-A FL Methods
IV-B Further Results and Discussions
  1. (Set 1)
  2. (Set 2)
These sets of experiments are discussed below in order.
Further, a major takeaway of implementing these technique were as follows: the tasks exhibit model bias at different numbers of local SGD iterations ($K$), which depends on the task complexity and the underlying data distributions. In other words, some task may need to be trained over longer duration of local training (i.e., higher values of $K$) to exhibit the model bias and subsequently a gap in performance under various levels of data heterogeneity. Nevertheless, to have fair comparisons between class-based and embedding-based methods, we maintained the same number of local SGDs per round for both class-based and embedding-based experiments across all benchmark methods and all tasks. The details of these hyperparameter are available at our aforementioned GitHub repository.

### v3

@@ Abstract
Federated Learning (FL) represents a paradigm shift in distributed machine learning (ML), enabling clients to train models collaboratively while keeping their raw data private. This paradigm shift from traditional centralized ML introduces challenges due to the non-iid (non-independent and identically distributed) nature of data across clients, significantly impacting FL’s performance. Existing literature, predominantly model data heterogeneity by imposing label distribution skew across clients. In this paper, we show that label distribution skew fails to fully capture the real-world data heterogeneity among clients in computer vision tasks beyond classification. Subsequently, we demonstrate that current approaches overestimate FL’s performance by relying on label/class distribution skew, exposing an overlooked gap in the literature. By utilizing pre-trained deep neural networks to extract task-specific data embeddings, we define task-specific data heterogeneity through the lens of each vision task and introduce a new level of data heterogeneity called embedding-based data heterogeneity. Our methodology involves clustering data points based on embeddings and distributing them among clients using the Dirichlet distribution. Through extensive experiments, we evaluate the performance of different FL methods under our revamped notion of data heterogeneity, introducing new benchmark performance measures to the literature. We further unveil a series of open research directions that can be pursued.
@@ I-A Overview and Literature Review
To bridge this gap, we redefine the notion of heterogeneity in FL for computer vision tasks as follows:
@@
To achieve the above objective, for the first time in the literature, we introduce a new layer of data heterogeneity, which we call embedding-based data heterogeneity. To model embedding-based data heterogeneity (i.e., embedding distribution skew), we leverage a pre-trained deep neural network (DNN) for each task and feed forward data points through this network. We then exploit the fact that the initial layers of the DNN of each task extract low-level features. As the data progresses through deeper layers, the network extracts higher-level representations that are increasingly specific to the vision task of interest. As we reach the layer just prior to the last layer (i.e, penultimate layer), embeddings that encapsulate a comprehensive view of the data from the perspective of the task are obtained, integrating both low-level and high-level features. We then leverage K-means to cluster datapoints based on these embeddings. Subsequently, we utilize a Dirichlet distribution to assign data points of each cluster to different clients to emulate task-specific data heterogeneity.
@@ I-B Summary of Contributions
We highlight the significance of task similarity in FL, illustrating that similar tasks tend to perceive data heterogeneity in a consistent manner, while differing tasks may interpret the same data distribution across clients with varying degrees of heterogeneity. Through this nuanced observation we unveil a series of novel research directions on addressing data heterogeneity for vision tasks in single-task and multi-task FL settings.
@@ II-B4 FL for Non-Classification Vision Tasks under Class-based Data Heterogeneity
We next take one of the well-known computer vision datasets, called Taskonomy [15], due to the large number of tasks it includes which is why it has been previously used in both centralized [23] and federated [16] multi-task learning as a sole baseline. We utilize the above-described Dirichlet distribution to create a heterogeneous allocation of data across $25$ clients’ local datasets in FL using the available class labels, referred to as “class_scene” in this dataset. [^4] The classes are provided in the dataset as probabilities for each datapoint belonging to each available class from which we choose the class with the highest probability as the determined class for each datapoint. More details on the data processing is provided in Appendix A. In our experiments, we select a representative set of vision tasks in this dataset, which are visualized in Fig. 2 and include “2D Edges,” “Reshading,” “Surface Normals,” “Semantic Segmentation,” “3D Keypoints,” “Euclidean Depth Estimation,” and “Scene Classification.”
These vision tasks – except Scene Classification – can be construed as non-classification tasks since they are associated with a class label. Following the non-iid data induced by Dirichlet distribution across the clients, we study the performance of the FL global model under different levels of heterogeneity – various $\alpha$ values – for the aforementioned vision tasks. The results of this experiment are shown in Fig. 3 (the left plot in each box, titled “Class-based”).
@@ III-B Emulating Heterogeneous Local Datasets for Clients
Formally, once the clustering of datapoints based on their embeddings are obtained, we emulate the local datasets $\mathcal{D}_{n}$ for each client $n$, where the set of ratios for classes $\mathbf{p}$ detailed in Sec. II-B3 is now replaced by the distribution of the embeddings in each cluster/group, which we denote by $\widehat{\mathbf{p}}$. Also, we let the set $\mathcal{G}$ contain the distinct cluster/group indices (e.g., all $16$ groups in Fig. 5). In our approach, each client $n$ will be assigned data from each cluster/group based on its own cluster/group distribution vector $\widehat{\mathbf{q}}_{n}=\mathsf{Dir}(\alpha\widehat{\mathbf{p}})=\{\widehat{q}_{n,g}\}_{g\in\mathcal{G}}$ where $\widehat{q}_{n,g}$ denotes the ratio of client $n$’s dataset containing data from group $g$ (e.g., one of the $16$ groups in Fig. 5). These local datasets are then used during the local training phase of FL.
@@
Remark 2.
@@
[Implications on the Concept of Task Similarity] Following the footsteps of prior work on modeling task similarity in the centralized ML domain [28, 29], our approach opens a model-driven perspective to quantify the similarity between two tasks. Specifically, this involves clustering data points based on their embeddings obtained from the models pre-trained for each task and comparing the resulting clusters to assess similarity. Using this straightforward approach, we establish a baseline performance for task similarity analysis, with results discussed in Sec. IV-B2. While this serves as a preliminary exploration, more sophisticated methods for task similarity measurement in the FL setting remain an open area for future investigation. Furthermore, this analysis provides an avenue for developing new task grouping strategies in FL, akin to prior efforts in centralized ML [30]. For example, by employing the task similarity measures based on the embedding-based clustering of data points, tasks with highly similar clustering patterns can be grouped and trained collaboratively, potentially enhancing FL performance for related tasks. We leave further exploration of this notion as future work.
Remark 3.
@@
[Implications on Multi-Task FL and its Extensions] A fundamental implications of our method lies in its application to multi-task FL and its extensions. In conventional multi-task FL [19, 18], each client typically has a single local task, with tasks varying across clients. On the other hand, the extended variations of multi-task FL [16, 17] consider scenarios where multiple tasks coexist on a single FL client. In the context of conventional multi-task FL, our method enables researchers to explore how varying levels of data heterogeneity, from the perspectives of different tasks, affect the performance of multi-task FL systems. This opens the door to designing countermeasures specifically aimed at addressing task-induced data heterogeneity in multi-task FL. Similarly, for the extensions of multi-task FL, our method allows researchers to investigate how inducing data heterogeneity from the perspective of one task impacts the performance of other tasks trained concurrently on the same client. Moreover, it facilitates the study of grouping and training local tasks together under different levels of data heterogeneity, potentially leading to new strategies for optimizing task interactions in FL. We leave further exploration of these concepts to future work.
IV Experiments and Benchmarks
IV-A FL Methods
  1. 1.
    FedAvg[2]: One of the most widely used methods in FL with the dynamics explained in Sec. II-A. The initial results presented in Fig. 3 were also gathered by employing this method.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
IV-B Further Results and Discussions
  1. (Set 1)
  2. (Set 2)
These sets of experiments are discussed below in order.
Furthermore, a major takeaway of implementing these technique were as follows: the tasks exhibit model bias at different numbers of local SGD iterations ($K$), which depends on the task complexity and the underlying data distributions. In other words, some task may need to be trained over longer duration of local training (i.e., higher values of $K$) so as to exhibit the model bias and subsequently a gap in performance under various levels of data heterogeneity. Nevertheless, to have fair comparisons between class-based and embedding-based methods, we maintained the same number of local SGDs per round for both class-based and embedding-based experiments across all benchmark methods and all tasks. The details of these hyperparameter are available at our aforementioned GitHub repository.